OpenStudy (anonymous):

please help someone anyone four and two thirds divided by one and one-sixth equals what value?

OpenStudy (anonymous):

4 and 2/3= 4+2/3=14/3 1 and 1/6=7/6 (14/3)/(7/6)= 14*6/(3*7)=2*2=4

OpenStudy (anonymous):

@gauravkj can you help me please Find the quotient: four-ninths divided by eight.

OpenStudy (anonymous):

1/18